OpenStudy (anonymous):

What are the steps to solve this problem? Factoring Quadratic Equations. 12(x)^(2)-41(x)+24=0

OpenStudy (mertsj):

Use the quadratic formula. a = 12, b = -41, c = 24

OpenStudy (anonymous):

Could you work it out for me? I have 12.24=288 . 32 and 9 are common den..

hero (hero):

You don't need the quadratic equation to solve this. You need to factor it

OpenStudy (mertsj):

\[x=\frac{41\pm \sqrt{41^2-4(12)(24)}}{2(12)}=\frac{41\pm \sqrt{1681-1152}}{24}\]

OpenStudy (mertsj):

Ok. Now the real brain will help you. Good bye.

OpenStudy (anonymous):

thank you

hero (hero):

12x^2-41x+24=0 12x^2 - 32x - 9x + 24 = 0 4x(3x - 8) - 3(3x - 8) = 0 (3x - 8)(4x - 3) = 0

OpenStudy (anonymous):

Thank you Hero. You are my Hero!!! I get it now!!!

hero (hero):

You figure it out by finding two numbers that multiply to get 12*24, yet add to get -41

hero (hero):

mn = 288 m + n = -41 The only two numbers that work are m = -32, n = -9

OpenStudy (anonymous):

Yes I had that right. not nothing else. Thanks a lot!!!